OpenStudy (anonymous):

one number is twice another number. If 12 is subtracted from the larger number the result is 7 more than the smaller number. find the number

OpenStudy (anonymous):

2x

OpenStudy (anonymous):

Let the No be x and 2x 2x - 12 = x+7 x = 19 2x = 38

OpenStudy (anonymous):

@Yahoo! i got 2x = 19 which makes x 9.5

OpenStudy (anonymous):

@xokatexo That's why you need to go step by step :P

OpenStudy (anonymous):

@Yahoo! i did 2x-12 = x+7 +12 +12 ____________________________ 2x = 19

OpenStudy (anonymous):

@Chlorophyll

OpenStudy (anonymous):

lol...u misse one x

OpenStudy (anonymous):

ohhh okay i got it

OpenStudy (anonymous):

@Chlorophyll @Yahoo! THANKS SO MUCH